OpenStudy (anonymous):

x = 4^{log_{4} 9 + log_{4} 4}

OpenStudy (anonymous):

\[\huge 4^{\log_4(9)+\log_4(4)}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

first off it is the same as \[4^{\log_4(36)}\]

OpenStudy (anonymous):

and since the log is the inverse of the exponential, you get \(36\) as a final answer

OpenStudy (anonymous):

you multiply it?

OpenStudy (anonymous):

\[\log(A)+\log(B)=\log(AB)\]

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

\frac{9}{7} \log_{5} x = -6

OpenStudy (anonymous):

what do i do? 9/7 log base 5 of x= -6

OpenStudy (anonymous):

\[\huge \frac{9}{7}\log_5(x)=-6\]?

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

multiply by \(\frac{7}{9}\) first

OpenStudy (anonymous):

then

OpenStudy (anonymous):

-4.6666

OpenStudy (anonymous):

start with \[\log_5(x)=-\frac{14}{3}\] then \[\large x=5^{-\frac{14}{3}}\] whatever that is

OpenStudy (anonymous):

oh you're very helpful thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

it wont come out as a real answer

myininaya (myininaya):

you can write that as 5^(-n) as 1/(5^n)

OpenStudy (anonymous):

still get a weird number

myininaya (myininaya):

well you can write it as \[\frac{1}{5^{\frac{14}{3}}}=\frac{1}{5^\frac{3(4)+2}{3}}=\frac{1}{ 5^4+5^\frac{2}{3}}\]

myininaya (myininaya):

but not sure that looks any better

OpenStudy (anonymous):

what counts as a weird number? i have heard of rational irrational transcendental but not "weird"

myininaya (myininaya):

all of those look weird and funny to me

myininaya (myininaya):

except rational i like rationals they are cute and cuddly

OpenStudy (anonymous):

help!!!!!!!

myininaya (myininaya):

@zencvbnmjkl are you doing a multiple choice thing or something and you don't that answer ?

myininaya (myininaya):

is that why you refer to it as weird?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it doesnt work in the calculator

OpenStudy (anonymous):

@satellite73

myininaya (myininaya):

it should work in the calculator

myininaya (myininaya):

ok type this:

myininaya (myininaya):

1 division sign ( 5 ^ (14/3) )

myininaya (myininaya):

make sure you put the parenthesis

OpenStudy (anonymous):

i got it apperciate it

myininaya (myininaya):

np